# Sourced Carrollian Fluids Dual to Black Hole Horizons

### v2

@@ Appendix D Details of numerical implementation
In the case for which $s_{1}+s_{2}+s_{3}=0$, these integrals can be computed via Wigner-$3j$ symbols as in [55]:
@@
Notably, for the expression we obtain, the condition $s_{1}+s_{2}+s_{3}=0$ is satisfied, and the numerical integration of the harmonics can be avoided using the $3j$-symbols in SymPy [80].
@@ D.1 Scalar field source profile
In the perturbed Schwarzschild model we study, we use Schwarzschild QNM frequencies for scalar modes [88, 89]. Specifically, we used the source modes $(\ell,m)=(0,0)$, $(1,1)$, (1,1)(1,-1), and $(2,0)$ as illustrated in Fig. 1 in the main text. For simplicity, we set $v_{0}=0$ and $\sigma=1$.
@@ E.2 Slowly-rotating Kerr
The slow rotation limit of Kerr spacetime is obtained from the standard Boyer–Lindquist coordinates by setting $a<<1$, where $a\equiv J/M$ [90]. The line element in this limit becomes:
@@
These new coordinates are motivated in [90] as they are constant on the generators of null hypersurfaces defined by $v=\mathrm{constant}$. Thus, we obtained the following metric:
@@
Finally, we compute the horizon angular momentum using the Hájíček field. As the spacetime is axisymmetric, there exists the axial Killing vector which can be expressed in these coordinates as $\partial_{{\psi}}$. Following Damour’s derivation [3], we compute the angular momentum of the horizon via
@@
The identifications from the slowly-rotating Kerr spacetime to the geometric quantities that are useful for the Carrollian fluid/horizon dynamics dictionary can be extended to the full Kerr solution at the cost of trading the simple results of this appendix for much more complex relations.

### v3

@@ Appendix D Details of numerical implementation
In the case for which $s_{1}+s_{2}+s_{3}=0$, these integrals can be computed via Wigner-$3j$ symbols as in [Redondo_Yuste_2023]:
@@
Notably, for the expression we obtain, the condition $s_{1}+s_{2}+s_{3}=0$ is satisfied, and the numerical integration of the harmonics can be avoided using the $3j$-symbols in SymPy [10.7717/peerj-cs.103].
@@ D.1 Scalar field source profile
In the perturbed Schwarzschild model we study, we use Schwarzschild QNM frequencies for scalar modes [Berti_2009, Stein:2019mop]. Specifically, we used the source modes $(\ell,m)=(0,0)$, $(1,1)$, (1,1)(1,-1), and $(2,0)$ as illustrated in Fig. 1 in the main text. For simplicity, we set $v_{0}=0$ and $\sigma=1$.
@@ E.2 Slowly-rotating Kerr
The slow rotation limit of Kerr spacetime is obtained from the standard Boyer–Lindquist coordinates by setting $a<<1$, where $a\equiv J/M$ [Poisson:2009pwt]. The line element in this limit becomes:
@@
These new coordinates are motivated in [Poisson:2009pwt] as they are constant on the generators of null hypersurfaces defined by $v=\mathrm{constant}$. Thus, we obtained the following metric:
@@
Finally, we compute the horizon angular momentum using the Hájíček field. As the spacetime is axisymmetric, there exists the axial Killing vector which can be expressed in these coordinates as $\partial_{{\psi}}$. Following Damour’s derivation [Damour1982], we compute the angular momentum of the horizon via
@@
The identifications from the slowly-rotating Kerr spacetime to the geometric quantities that are useful for the Carrollian fluid/horizon dictionary can be extended to the full Kerr solution at the cost of trading the simple results of this appendix for much more complex relations.